OpenStudy (anonymous):

Find the perimeter of the rectangle with 13ft as one side and 8ft as the other.

OpenStudy (anonymous):

42

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

13x8?

OpenStudy (anonymous):

the perimeter is 2L + 2W = 2(13) + 2(8) = 26 + 16 = 42

OpenStudy (anonymous):

13*8 would be the area.. not the perimeter